niel is three years older than his sister together there ages add up to 27 how old are they
Answer by Cromlix(4381)

Hi there,
Make neil's sister = x
neil is x + 3
x + (x + 3) = 27
x + x + 3 = 27
Collect like terms.
x + x = 27 - 3
2x = 24
x = 12.
Neil is 15 years old
Neil's sister is 12 years old.
Hope this helps :-)